When Caroline's dress factory hires two workers, the total product is 50 dresses. When she hires three workers, the total product is 60, and when she hires four workers, the total product is 65. The slope of the marginal product curve between two and four workers hired is:

A.infinite.

B. negative.

C. positive.

D. zero.
Answer #1

ANswer

negative


MP(n)=(TP(n)-TP(p))/(n-p)
MP(n)=marginal product of n th unit of input
TP(n)= total product of n inputs
TP(p)=total product of p inputs
it is true for n>p

MP(3)=(60-50)/(3-2)=10

MP(4)=(65-60)/(4-3)=5

the marginal product is decreasing means increasing in labor and marginal product has negative slope.

TP 1 2 3 4 5 6 7 89 10 Labor 1 2 3 45 6 7 9 10 Labor MP